# This Time is Different: An Observability Perspective on Time Series Foundation Models

### v1

@@ A.2.1 Complexity analysis
In the time-wise attention blocks, we use causal masking and rotary positional embeddings [43] with XPOS [44] in order to autoregressively model time-dependent features.
@@ A.3 Probabilistic prediction
Practitioners who rely on time series forecasting typically prefer probabilistic predictions. A common practice in neural time series models is to use an output layer where the model regresses the parameters of a probability distribution. This allows for prediction intervals to be computed using Monte Carlo sampling (see Appendix A.4) [84].
Common choices for an output layer are Normal [84] and Student-T [85, 21], which can improve robustness to outliers. Moirai [13] allows for more flexible residual distributions by proposing a novel mixture model incorporating a weighted combination of Gaussian, Student-T, Log-Normal, and Negative-Binomial outputs.
However, real-world time series can often have complex distributions that are challenging to fit, with outliers, heavy tails, extreme skew, and multimodality. In order to accommodate these scenarios, we introduce an even more flexible output likelihood in Section 3.1 based on a Student-T mixture model [54].
@@
In our ablation study (Appendix E), we find that the SMM improves both point prediction and probabilistic forecasting accuracy when compared with a single Student-T distribution as used in TiDE [85], Lag-Llama [21], and implementations of DeepAR [84], PatchTST [19], iTransformer [42], and others in the popular open-source GluonTS library [86].
@@
where each equation produces a vector in $\mathbb{R}^{K}$ containing the parameters for all mixture components at time $t$. The individual component parameters $\nu_{t,k}$, $\mu_{t,k}$, $\tau_{t,k}$, and $\tilde{\pi}_{t,k}$ (the mixture logits) are the $k$th elements of these vectors. The parameter $\epsilon$ is machine epsilon (the smallest positive floating-point number), and softplus($x$) = $\log(1+e^{x})$. The use of softplus and $\epsilon$ ensure that the scale $\tau$ remains positive. Similarly, we add the constraint $\nu>2$ to ensure that each component of our mixture has well-defined first and second moments (mean and variance).
@@ A.4 Forecasting
When performing inference, we draw $u$ (for some user specified integer $u>0$) samples from the mixture distribution at each timestamp, then feed each sample back into the decoder for the next prediction, resulting in $n$ identically and independently sampled time-series. This allows us to produce prediction intervals at any quantile, limited only by the number of samples. Our exact sampling procedure for several tasks is detailed in Section 5.
@@ A.5 Loss function
For a ground truth value $x_{t}$ in patch $i+1$ and the mean prediction $\hat{x}_{t}=\mathbb{E}[p(x_{t}|X_{1:i})]$, the robust loss is defined below [60]:
@@
Here, $\mathcal{L}_{\text{Robust}(\alpha,\delta)}$ serves as a point prediction error measure, where $\alpha\leq 2$ is a shape parameter that controls the robustness to outlier observations (Fig. 7) and $\delta>0$ is a scale parameter that determines the size of the parabolic portion of the loss curve. This loss component directly penalizes point prediction accuracy, and we conjecture this may help steer the mixture model away from degenerate solutions of the type described in [59]. In our ablation study, we find that adding the robust loss component significantly improves point forecasting accuracy without hurting probabilistic predictions (Section E).
@@ A.6 Hyperparameter optimization
To determine the optimal architecture and training configuration for Toto, we conducted an extensive hyperparameter sweep using Optuna [87], a Bayesian optimization framework. We employed the Tree-structured Parzen Estimator (TPE) algorithm with 65 trials to efficiently explore the high-dimensional search space.
Our optimization objective was to minimize the validation mean absolute error (MAE) on multi-step forecasting tasks on a random validation split of the observability portion of the pretraining data. We train the model using the AdamW optimizer [88] with a WSD learning rate scheduler [89]. We performed this sweep over 50,000 steps over 133 iterations over ranges described in Table 5.
@@
Learning Rate $[10^{-5},5\times 10^{-3}]$
Weight Decay $[10^{-3},10^{-1}]$
@@
The resulting hyperparameter configuration described in Table 6 obtained the best multistep (average of 96 and 192) MAE on the observability validation set.
@@
In Section E, we perform an ablation study on the impact of various model components. We optimize speed and memory usage by utilizing fused kernel implementations and memory efficient attention operations via xformers [90], (with the FlashAttention-3 kernel [91]).
@@ B.1 Observability dataset
Figure 8: Example metric query in the observability platform. The metric name (1) determines which metric is being queried. The filter clause (2) limits which contexts are queried, in this case restricting the query to apps in the prod environment. The space aggregation (3) indicates that the sum of the metric value should be returned for each unique value of the group-by key(s), aggregated across all other keys. The time aggregation (4) indicates that metric values should be aggregated to the average for each 60-second interval. The query results will be a multivariate time series with 1-minute time steps, and with separate individual variates for each unique value of cluster_name.
@@ B.2 Public datasets
We train on a public dataset corpus, which exposes the model to diverse time series behaviors across different domains and sampling frequencies. Our pre-training dataset incorporates a diverse collection of time series from the GIFT-Eval Pretrain collection [26] and non-overlapping Chronos datasets [11].
@@ B.3 Synthetic data
We supplement our training with synthetic data to further improve model performance. Our synthetic dataset consists of procedurally generated time series using an approach similar to TimesFM [12], as well as kernel_synth_1m from the Chronos dataset [11]. Synthetic data constitutes approximately 33% of our training dataset.
@@ B.4 Preprocessing
Next, various data augmentations are employed to enhance the dataset's robustness. We introduce random time offsets to prevent memorization caused by having series always align the same way with the patch grid. After concatenating the observability and public datasets for training, we also implement a variate shuffling strategy to maintain diversity and representation. Specifically, we randomly combine variates from either observability, open source datasets (GIFT-Eval pretrain and Chronos datasets), and/or synthetic data with a probability of  14%, thus creating new, diverse combinations of data points. We shuffle series with adjacent indices (batched by 32 variates), favoring data points that were closer together in the original datasets. This approach improves the model's ability to generalize across different types of data effectively.
@@ C.1 Domain taxonomy
Figure 10: Representative examples from Boom, illustrating the unique temporal patterns associated observability data.
Figure 11: 2D Principal Component Analysis [92] projection of normalized statistical features computed from the observability Benchmark and GIFT-Eval datasets. The clear separation between the two distributions highlights a significant shift in underlying time series characteristics.
@@ C.2.2 Evaluation metrics
MASE [76], a point-forecasting score, is defined as
@@
CRPS [77] is scoring rule for probabilistic forecasts, defined with respect to an observation $y$
[…]
where $F_{D}$ is the cumulative distribution function of the forecast distribution $D$ and $H$ is the Heaviside step function. We follow the standard practice of taking the mean weighted quantile loss as a discrete approximation, as described by Park et al. [78].
@@ C.2.3 Inference procedures
To evaluate the comparison models on BOOM, we closely follow the evaluation methodology used in the GIFT-Eval implementation. For models not included in GIFT-Eval, we rely on their official implementations and recommended evaluation procedures. All foundation models are evaluated using a unified context length of 2048. This choice is informed by preliminary experiments showing that a shorter context length (512) leads to a general degradation in performance across models. Therefore, we opt for a relatively large context window (2048) to preserve forecast quality, while ensuring feasibility on available hardware.
@@
For the three statistical baselines—AutoARIMA, AutoTheta, and AutoETS—we use the default hyperparameter settings from the statsforecast package, with one exception: for AutoARIMA, we reduce $max_{d}$ and $max_{D}$ from 2 to 1 due to frequent numerical instability when d = D = 2. Following GIFT-Eval, we set the maximum input length for all statistical models to 1000.
@@ C.2.4 Aggregation of results
Common practice for aggregating normalized benchmarking statistics is to use the geometric mean [93]. We adopt this approach, with a slight caveat: due to the presence of constant subsequences in observability time series, it’s possible for models to achieve zero error on a handful of series in Boom. As zeros cause geometric means to collapse, we instead use the shifted geometric mean, which is stable in such scenarios [79, 80], for aggregating MASE and CRPS. Specifically, for a metric $m$ across a set of test instances $N$, the shifted geometric mean $\bar{m}_{\text{ShiftedGeom}}$ is defined as
@@
NaN values.: The benchmark includes a small number of zero-inflated series, which can result in invalid CRPS values due to division-by-zero errors. To mitigate this, we apply a generic postprocessing strategy where NaN and infinite values are imputed using the mean CRPS across the remaining series.
Extremely low naive errors.: Certain flat series exhibit extremely low or even null in-sample MAE, leading to instability when normalizing by the seasonal naive MAE, with potentially exploding normalized values. Rather than discarding these special cases — which are informative for evaluating how models handle anomalous flat behavior — we isolate them into a separate data split. This split is evaluated using simple MAE instead of MASE and non-normalized CRPS. To define this subset, we apply an objective criterion: we include all series where either the MAE for the Seasonal Naive forecast on the test set or the in-sample Seasonal Naive MAE 13 are zero. If any prediction window of any variate within a query satisfies the above conditions, we assign the entire query to the second, low-variability set, where metrics are computed without normalization. For evaluation, results in this subset are not normalized and thus are aggregated using a simple arithmetic mean.
@@
Taxonomy breakdowns.. For the Boom, we evaluate model performance across stratified groups defined by the dataset’s taxonomy categories. For term (Table 12), metric type (Table 13), and domain (Table 14), we report the shifted geometric mean of the evaluation metric within each group. These results are discussed in detail in Section D.1.
@@ C.3 Boomlet
To facilitate research in settings where it is infeasible to evaluate on the full Boom benchmark, we construct a smaller, representative subset, denoted as Boomlet. This subset is obtained via uniform sampling over metric queries in Boom, ensuring preservation of the distributional properties across key taxonomy dimensions. Boomlet comprises 32 metric queries, encompassing a total of 1,627 variates and approximately 23 million observation points.
@@ D.1 BOOM
In Fig. 12, we present qualitative comparisons across three representative forecasting scenarios to highlight the behavioral differences between Toto, Chronos, and Moirai. In the first example (a), features a highly stochastic signal interwoven with complex seasonal components. While Moirai and Chronos models tend to overfit short-term fluctuations—resulting in jagged forecasts and unstable confidence intervals— Toto effectively identifies and extrapolates the latent seasonal structure, yielding smoother, more coherent trajectories and uncertainty bands that reflect a deeper structural understanding of the series dynamics. Example (b) the target signal exhibits high dynamism with rapidly oscillating structure and sustained amplitude modulations—posing a challenge for long-range temporal modeling. While both Moirai and Chronos models progressively lose phase alignment and dampen their amplitude estimates, Toto consistently maintains sharp, temporally aligned forecasts with well-calibrated uncertainty, accurately tracking the intricate periodic structure far into the forecast horizon. Finally, example (c), the target series is characterized by sparse, bursty impulses with high variance across events. Here, although Toto ’s mean prediction does not always precisely capture individual peaks, its predictive distribution faithfully mirrors the underlying spikiness of the series, in stark contrast to Chronos, which collapses to an overconfident flat trajectory.
@@
Dataset Metric Toto MoiraiSmall MoiraiBase MoiraiLarge TimesFM2.0 Chronos-BoltSmall Chronos-BoltBase Timer Time-MoELarge Time-MoEBase VisionTS Naive
Boom MASE $\downarrow$ 0.617 $0.738$ 0.710 $0.720$ $0.725$ $0.733$ $0.726$ $0.796$ $0.881$ $0.806$ $0.991$ $1.000$
CRPS $\downarrow$ 0.375 $0.447$ 0.428 $0.436$ $0.447$ $0.455$ $0.451$ $0.639$ $0.643$ $0.649$ $0.675$ $1.000$
Boomlet MASE $\downarrow$ 0.617 $0.786$ 0.779 $0.767$ 0.685 $0.717$ $0.711$ $0.807$ $0.793$ $0.810$ $0.912$ $1.000$
CRPS $\downarrow$ 0.519 $0.631$ 0.630 $0.621$ 0.603 $0.642$ $0.637$ $0.793$ $0.780$ $0.788$ $0.885$ $1.000$
@@
Term Metric Toto MoiraiSmall MoiraiBase MoiraiLarge TimesFM2.0 Chronos-boltSmall Chronos-boltBase Timer Time-MoEBase VisionTS Naive
Long MASE $\downarrow$ 0.688 $0.795$ 0.780 $0.799$ $0.817$ $0.813$ $0.798$ $0.809$ $0.886$ $1.026$ $1.000$
CRPS $\downarrow$ 0.424 $0.482$ 0.473 $0.491$ $0.522$ $0.528$ $0.519$ $0.661$ $0.724$ $0.698$ $1.000$
Medium MASE $\downarrow$ 0.657 $0.771$ 0.753 $0.770$ $0.780$ $0.782$ $0.782$ $0.804$ $0.866$ $1.011$ $1.000$
CRPS $\downarrow$ 0.406 $0.476$ 0.460 $0.475$ $0.499$ $0.508$ $0.507$ $0.671$ $0.725$ $0.698$ $1.000$
Short MASE $\downarrow$ 0.535 $0.670$ $0.627$ $0.626$ 0.619 $0.638$ $0.632$ $0.779$ $0.704$ $0.947$ $1.000$
CRPS $\downarrow$ 0.318 $0.399$ $0.370$ $0.369$ 0.359 $0.368$ $0.365$ $0.597$ $0.541$ $0.640$ $1.000$
@@
Length Metric Toto MoiraiSmall MoiraiBase MoiraiLarge TimesFM2.0 Chronos-BoltSmall Chronos-BoltBase Timer Time-MoEBase VisionTS Naive
Count MASE $\downarrow$ $0.687$ $0.814$ $0.795$ $0.813$ $0.919$ $0.883$ $0.880$ 0.663 0.652 $1.220$ $1.000$
CRPS $\downarrow$ 0.317 $0.370$ 0.353 $0.372$ $0.403$ $0.403$ $0.402$ $0.662$ $0.651$ $0.603$ $1.000$
Distribution MASE $\downarrow$ 0.658 $0.741$ 0.724 $0.729$ $0.745$ $0.759$ $0.753$ $0.890$ $0.878$ $1.034$ $1.000$
CRPS $\downarrow$ 0.382 $0.434$ 0.422 $0.428$ $0.440$ $0.452$ $0.446$ $0.608$ $0.604$ $0.674$ $1.000$
Gauge MASE $\downarrow$ 0.583 $0.720$ 0.686 $0.700$ $0.706$ $0.706$ $0.696$ $0.721$ $0.760$ $0.922$ $1.000$
CRPS $\downarrow$ 0.382 $0.471$ 0.444 $0.456$ $0.466$ $0.469$ $0.463$ $0.658$ $0.694$ $0.672$ $1.000$
Rate MASE $\downarrow$ 0.634 $0.753$ $0.728$ $0.733$ 0.726 $0.742$ $0.739$ $0.864$ $0.846$ $1.041$ $1.000$
CRPS $\downarrow$ 0.369 $0.433$ 0.418 $0.422$ $0.431$ $0.445$ $0.443$ $0.630$ $0.619$ $0.687$ $1.000$
Table 13: [Metric type] Performance (MASE and CRPS) of TOTO and other zero-shot models on the BOOM dataset across different horizon lengths. Key: Best results, Second-best results.
Domain Metric Toto MoiraiSmall MoiraiBase MoiraiLarge TimesFM2.0 Chronos-BoltSmall Chronos-BoltBase Timer Time-MoEBase VisionTS Naive
Application usage MASE $\downarrow$ 0.639 $0.747$ 0.721 $0.730$ $0.736$ $0.748$ $0.748$ $0.871$ $0.863$ $1.042$ $1.000$
CRPS $\downarrow$ 0.378 $0.440$ 0.422 $0.430$ $0.441$ $0.452$ $0.451$ $0.636$ $0.633$ $0.691$ $1.000$
Database MASE $\downarrow$ 0.635 $0.751$ $0.738$ $0.743$ $0.765$ $0.761$ $0.757$ $0.716$ 0.714 $1.017$ $1.000$
CRPS $\downarrow$ 0.362 $0.429$ 0.414 $0.418$ $0.440$ $0.444$ $0.441$ $0.619$ $0.618$ $0.647$ $1.000$
Infrastructure MASE $\downarrow$ 0.568 $0.692$ 0.650 $0.670$ $0.679$ $0.678$ $0.663$ $0.728$ $0.791$ $0.863$ $1.000$
CRPS $\downarrow$ 0.391 $0.476$ 0.446 $0.462$ $0.471$ $0.474$ $0.466$ $0.655$ $0.713$ $0.666$ $1.000$
Networking MASE $\downarrow$ 0.635 $0.795$ $0.786$ $0.773$ $0.765$ $0.779$ 0.757 $0.871$ $0.856$ $1.035$ $1.000$
CRPS $\downarrow$ 0.400 $0.493$ 0.484 0.484 $0.493$ $0.506$ $0.489$ $0.725$ $0.721$ $0.734$ $1.000$
Security MASE $\downarrow$ 0.682 $0.741$ $0.739$ $0.736$ 0.717 $0.734$ $0.729$ $0.828$ $0.770$ $0.924$ $1.000$
CRPS $\downarrow$ 0.476 $0.505$ 0.504 0.504 $0.525$ $0.539$ $0.535$ $0.664$ $0.625$ $0.735$ $1.000$

### v2

@@ A.2.1 Complexity analysis
In the time-wise attention blocks, we use causal masking and rotary positional embeddings [Su2024] with XPOS [sun2022a] in order to autoregressively model time-dependent features.
@@ A.3 Probabilistic prediction
Practitioners who rely on time series forecasting typically prefer probabilistic predictions. A common practice in neural time series models is to use an output layer where the model regresses the parameters of a probability distribution. This allows for prediction intervals to be computed using Monte Carlo sampling (see Appendix A.4) [Salinas2020].
Common choices for an output layer are Normal [Salinas2020] and Student-T [das2023longterm, rasul2023lagllama], which can improve robustness to outliers. Moirai [Woo2024] allows for more flexible residual distributions by proposing a novel mixture model incorporating a weighted combination of Gaussian, Student-T, Log-Normal, and Negative-Binomial outputs.
However, real-world time series can often have complex distributions that are challenging to fit, with outliers, heavy tails, extreme skew, and multimodality. In order to accommodate these scenarios, we introduce an even more flexible output likelihood in Section 3.1 based on a SMM [Peel2000].
@@
In our ablation study (Appendix E), we find that the SMM improves both point prediction and probabilistic forecasting accuracy when compared with a single Student-T distribution as used in TiDE [das2023longterm], Lag-Llama [rasul2023lagllama], and implementations of DeepAR [Salinas2020], PatchTST [Nie2023], iTransformer [shazeer2020gluvariantsimprovetransformer], and others in the popular open-source GluonTS library [gluonts_jmlr].
@@
where each equation produces a vector in $\mathbb{R}^{K}$ containing the parameters for all mixture components at time $t$. The individual component parameters $\nu_{t,k}$, $\mu_{t,k}$, $\tau_{t,k}$, and $\tilde{\pi}_{t,k}$ (the mixture logits) are the $k$th elements of these vectors. The parameter $\epsilon$ is a small positive constant, and softplus($x$) = $\log(1+e^{x})$. The use of softplus and $\epsilon$ ensure that the scale $\tau$ remains positive. Similarly, we add the constraint $\nu>2$ to ensure that each component of our mixture has well-defined first and second moments (mean and variance).
@@ A.4 Forecasting
When performing inference, we draw $u$ (for some user specified integer $u>0$) samples from the mixture distribution at each timestamp, then feed each sample back into the decoder for the next prediction, resulting in $n$ identically and independently sampled time-series. This allows us to produce prediction intervals at any quantile, limited only by the number of samples. Our exact sampling procedure for several tasks is detailed in Section C.2.
@@ A.5 Loss function
For a ground truth value $x_{t}$ in patch $i+1$ and the mean prediction $\hat{x}_{t}=\mathbb{E}[p(x_{t}|X_{1:i})]$, the robust loss is defined below [DBLP:journals/corr/Barron17]:
@@
Here, $\mathcal{L}_{\text{Robust}(\alpha,\delta)}$ serves as a point prediction error measure, where $\alpha\leq 2$ is a shape parameter that controls the robustness to outlier observations (Fig. 7) and $\delta>0$ is a scale parameter that determines the size of the parabolic portion of the loss curve. This loss component directly penalizes point prediction accuracy, and we conjecture this may help steer the mixture model away from degenerate solutions of the type described in [eisner_gmm_optimization]. In our ablation study, we find that adding the robust loss component significantly improves point forecasting accuracy without hurting probabilistic predictions (Section E).
@@ A.6 Hyperparameter optimization
To determine the optimal architecture and training configuration for Toto, we conducted an extensive hyperparameter sweep using Optuna [DBLP:journals/corr/abs-1907-10902], a Bayesian optimization framework. We employed the Tree-structured Parzen Estimator (TPE) algorithm to efficiently explore the high-dimensional search space.
Our optimization objective was to minimize the validation mean absolute error (MAE) on multi-step forecasting tasks on a random validation split of the observability portion of the pretraining data. We train the model using the AdamW optimizer [loshchilov2018decoupled] with a WSD learning rate scheduler [hu2024minicpmunveilingpotentialsmall]. We performed this sweep for 133 iterations over ranges described in Table 5, each for 50,000 training steps.
@@
Learning Rate $10^{-5},5\times 10^{-3}]$
Weight Decay $10^{-3},10^{-1}]$
@@
The resulting hyperparameter configuration described in Table 6 obtained the best multistep (average of 96 and 192) MAE on the Datadog validation set.
@@
In Section E, we perform an ablation study on the impact of various model components. We optimize speed and memory usage by utilizing fused kernel implementations and memory efficient attention operations via xformers [xFormers2022], (with the FlashAttention-3 kernel [shah2024flashattention3fastaccurateattention]).
@@ B.1 Observability dataset
Figure 8: Example metric query in the Datadog platform. The metric name (1) determines which metric is being queried. The filter clause (2) limits which contexts are queried, in this case restricting the query to apps in the prod environment. The space aggregation (3) indicates that the sum of the metric value should be returned for each unique value of the group-by key(s), aggregated across all other keys. The time aggregation (4) indicates that metric values should be aggregated to the average for each 60-second interval. The query results will be a multivariate time series with 1-minute time steps, and with separate individual variates for each unique value of cluster_name.
@@ B.2 Public datasets
We train on a public dataset corpus, which exposes the model to diverse time series behaviors across different domains and sampling frequencies. Our pre-training dataset incorporates a diverse collection of time series from the GIFT-Eval Pretrain collection [godahewa2021monash] and non-overlapping Chronos datasets [ansari2024chronoslearninglanguagetime].
@@ B.3 Synthetic data
We supplement our training with synthetic data to further improve model performance. Our synthetic dataset consists of procedurally generated time series using an approach similar to TimesFM [das2024a], as well as kernel_synth_1m from the Chronos dataset [ansari2024chronoslearninglanguagetime]. Synthetic data constitutes approximately 33% of our training dataset.
@@ B.4 Preprocessing
Next, various data augmentations are employed to enhance the dataset's robustness. We introduce random time offsets to prevent memorization caused by having series always align the same way with the patch grid. After concatenating the Datadog and public datasets for training, we also implement a variate shuffling strategy to maintain diversity and representation. Specifically, we randomly combine variates from either Datadog, open source datasets (GIFT-Eval pretrain and Chronos datasets), and/or synthetic data with a probability of  14%, thus creating new, diverse combinations of data points. We shuffle series with adjacent indices (batched by 32 variates), favoring data points that were closer together in the original datasets. This approach improves the model's ability to generalize across different types of data effectively.
@@ C.1 Domain taxonomy
Figure 10: Representative examples from the Boom, illustrating the unique temporal patterns associated observability data.
Figure 11: 2D Principal Component Analysis [Pearson01111901] projection of normalized statistical features computed from the Datadog Benchmark and GIFT-Eval datasets. The clear separation between the two distributions highlights a significant shift in underlying time series characteristics.
@@ C.2.2 Evaluation metrics
MASE [Hyndman2006], a point-forecasting score, is defined as
@@
CRPS [GneitlingCRPS] is scoring rule for probabilistic forecasts, defined with respect to an observation $y$
$\text{CRPS}(D,y)=\int_{\mathbb{R}}(F_{D}(x)-H(x-y))^{2}dx$ (14)
where $F_{D}$ is the cumulative distribution function of the forecast distribution $D$ and $H$ is the Heaviside step function. We follow the standard practice of taking the mean weighted quantile loss as a discrete approximation, as described by parkCRPS.
@@ C.2.3 Inference procedures
To evaluate the comparison models on Boom, we closely follow the evaluation methodology used in the GIFT-Eval implementation. For models not included in GIFT-Eval, we rely on their official implementations and recommended evaluation procedures.
All foundation models are evaluated using a unified context length of 2048. This choice is informed by preliminary experiments showing that a shorter context length (512) leads to a general degradation in performance across models. Therefore, we opt for a relatively large context window (2048) to preserve forecast quality, while ensuring feasibility on available hardware.
@@
For the three statistical baselines—AutoARIMA, AutoTheta, and AutoETS—we use the default hyperparameter settings from the statsforecast package, with one exception: for AutoARIMA, we reduce $max_{d}$ and $max_{D}$ from 2 to 1 due to frequent numerical instability when $d=D=2$.
Following GIFT-Eval, we set the maximum input length for all statistical models to 1000.
@@ C.2.4 Aggregation of results
Common practice for aggregating normalized benchmarking statistics is to use the geometric mean [Fleming1986HowNT]. We adopt this approach, with a slight caveat: due to the presence of constant subsequences in observability time series, it’s possible for models to achieve zero error on a handful of series in Boom. As zeros cause geometric means to collapse, we instead use the shifted geometric mean, which is stable in such scenarios [RavanaGmap, Beach2024], for aggregating MASE and CRPS. Specifically, for a metric $m$ across a set of test instances $N$, the shifted geometric mean $\bar{m}_{\text{ShiftedGeom}}$ is defined as
@@
NaN values: The benchmark includes a small number of zero-inflated series, which can result in invalid CRPS values due to division-by-zero errors. To mitigate this, we apply a generic postprocessing strategy where NaN and infinite values are imputed using the mean CRPS across the remaining series.
Extremely low naive errors: Certain flat series exhibit extremely low or even null in-sample MAE, leading to instability when normalizing by the seasonal naive MAE, with potentially exploding normalized values. Rather than discarding these special cases — which are informative for evaluating how models handle anomalous flat behavior — we isolate them into a separate data split. This split is evaluated using simple MAE instead of MASE and non-normalized CRPS. To define this subset, we apply an objective criterion: we include all series where either the MAE for the Seasonal Naive forecast on the test set or the in-sample Seasonal Naive MAE 13 are zero. If any prediction window of any variate within a query satisfies the above conditions, we assign the entire query to the second, low-variability set, where metrics are computed without normalization. For evaluation, results in this subset are not normalized and thus are aggregated using a simple arithmetic mean.
@@
Taxonomy breakdowns. For the Boom, we evaluate model performance across stratified groups defined by the dataset’s taxonomy categories. For term (Table 12), metric type (Table 13), and domain (Table 14), we report the shifted geometric mean of the evaluation metric within each group. These results are discussed in detail in Section D.1.
@@ C.3 Boomlet
To facilitate research in settings where it is computationally infeasible to evaluate on the full Boom benchmark, we construct a smaller, representative subset denoted as Boomlet. This subset is created via uniform sampling over metric queries in Boom, while additionally prioritizing queries with a relatively large number of variates to ensure sufficient training signal. This selection strategy preserves the distributional properties across key taxonomy dimensions, while also maintaining a practical volume of training data per query. Boomlet comprises 32 metric queries, encompassing 1,627 variates and approximately 23 million observation points.
@@ D.1 BOOM
[…]
@@
Dataset Metric Toto MoiraiSmall MoiraiBase MoiraiLarge TimesFM2.0 Chronos-BoltSmall Chronos-BoltBase Timer Time-MoE50M Time-MoE200M VisionTS DLinear DeepAR Naive
Boom MASE $\downarrow$ 0.617 $0.729$ 0.710 $0.720$ $0.725$ $0.733$ $0.726$ $0.796$ $0.806$ $0.881$ $0.988$ - - $1.000$
CRPS $\downarrow$ 0.375 $0.442$ 0.428 $0.436$ $0.447$ $0.455$ $0.451$ $0.639$ $0.649$ $0.643$ $0.673$ - - $1.000$
Rank $\downarrow$ 2.369 $4.905$ 4.328 $4.561$ $5.243$ $5.927$ $5.576$ $9.920$ $9.877$ $9.843$ $10.989$ - - $12.631$
Boomlet MASE $\downarrow$ 0.617 $0.786$ 0.779 $0.767$ 0.685 $0.717$ $0.711$ $0.807$ $0.810$ $0.793$ $0.912$ 0.823 0.883 $1.000$
CRPS $\downarrow$ 0.519 $0.631$ 0.630 $0.621$ 0.603 $0.642$ $0.637$ $0.793$ $0.788$ $0.780$ $0.885$ 0.641 0.697 $1.000$
Rank $\downarrow$ 1.300 $5.711$ $5.300$ $4.967$ 4.867 $6.756$ $6.511$ $11.544$ $11.222$ $11.189$ $13.589$ 6.056 7.900 $14.133$
@@
Real Term Metric Toto MoiraiSmall MoiraiBase MoiraiLarge TimesFM2.0 Chronos-BoltSmall Chronos-BoltBase Timer Time-MoEBase Time-MoELarge VisionTS Naive
Long MASE $\downarrow$ 0.688 $0.795$ 0.780 $0.799$ $0.817$ $0.813$ $0.798$ $0.809$ $0.886$ $0.950$ $1.026$ $1.000$
CRPS $\downarrow$ 0.424 $0.482$ 0.473 $0.491$ $0.522$ $0.528$ $0.519$ $0.661$ $0.724$ $0.694$ $0.698$ $1.000$
Medium MASE $\downarrow$ 0.657 $0.771$ 0.753 $0.770$ $0.780$ $0.782$ $0.782$ $0.804$ $0.866$ $0.929$ $1.011$ $1.000$
CRPS $\downarrow$ 0.406 $0.476$ 0.460 $0.475$ $0.499$ $0.508$ $0.507$ $0.671$ $0.725$ $0.692$ $0.698$ $1.000$
Short MASE $\downarrow$ 0.535 $0.670$ $0.627$ $0.626$ 0.619 $0.638$ $0.632$ $0.779$ $0.704$ $0.794$ $0.947$ $1.000$
CRPS $\downarrow$ 0.318 $0.399$ $0.370$ $0.369$ 0.359 $0.368$ $0.365$ $0.597$ $0.541$ $0.570$ $0.640$ $1.000$
@@
Type Metric Toto MoiraiSmall MoiraiBase MoiraiLarge TimesFM2.0 Chronos-BoltSmall Chronos-BoltBase Timer Time-MoEBase Time-MoELarge VisionTS Naive
Count MASE $\downarrow$ $0.687$ $0.814$ $0.795$ $0.813$ $0.919$ $0.883$ $0.880$ 0.663 0.652 $1.035$ $1.220$ $1.000$
CRPS $\downarrow$ 0.317 $0.370$ 0.353 $0.372$ $0.403$ $0.403$ $0.402$ $0.662$ $0.651$ $0.698$ $0.603$ $1.000$
Distribution MASE $\downarrow$ 0.658 $0.741$ 0.724 $0.729$ $0.745$ $0.759$ $0.753$ $0.890$ $0.878$ $0.877$ $1.034$ $1.000$
CRPS $\downarrow$ 0.382 $0.434$ 0.422 $0.428$ $0.440$ $0.452$ $0.446$ $0.608$ $0.604$ $0.596$ $0.674$ $1.000$
Gauge MASE $\downarrow$ 0.583 $0.720$ 0.686 $0.700$ $0.706$ $0.706$ $0.696$ $0.721$ $0.760$ $0.890$ $0.922$ $1.000$
CRPS $\downarrow$ 0.382 $0.471$ 0.444 $0.456$ $0.466$ $0.469$ $0.463$ $0.658$ $0.694$ $0.703$ $0.672$ $1.000$
Rate MASE $\downarrow$ 0.634 $0.753$ $0.728$ $0.733$ 0.726 $0.742$ $0.739$ $0.864$ $0.846$ $0.862$ $1.041$ $1.000$
CRPS $\downarrow$ 0.369 $0.433$ 0.418 $0.422$ $0.431$ $0.445$ $0.443$ $0.630$ $0.619$ $0.596$ $0.687$ $1.000$
Table 13: Performance comparison of Toto and other zero-shot models across different metric types. MASE and CRPS are normalized by the Seasonal Naive forecast and aggregated across tasks using the shifted geometric mean. Key: Best results, Second-best results.
Domain Metric Toto MoiraiSmall MoiraiBase MoiraiLarge TimesFM2.0 Chronos-BoltSmall Chronos-BoltBase Timer Time-MoEBase Time-MoELarge VisionTS Naive
Application usage MASE $\downarrow$ 0.639 $0.747$ 0.721 $0.730$ $0.736$ $0.748$ $0.748$ $0.871$ $0.863$ $0.884$ $1.042$ $1.000$
CRPS $\downarrow$ 0.378 $0.440$ 0.422 $0.430$ $0.441$ $0.452$ $0.451$ $0.636$ $0.633$ $0.611$ $0.691$ $1.000$
Database MASE $\downarrow$ 0.635 $0.751$ $0.738$ $0.743$ $0.765$ $0.761$ $0.757$ $0.716$ 0.714 $0.903$ $1.017$ $1.000$
CRPS $\downarrow$ 0.362 $0.429$ 0.414 $0.418$ $0.440$ $0.444$ $0.441$ $0.619$ $0.618$ $0.633$ $0.647$ $1.000$
Infrastructure MASE $\downarrow$ 0.568 $0.692$ 0.650 $0.670$ $0.679$ $0.678$ $0.663$ $0.728$ $0.791$ $0.847$ $0.863$ $1.000$
CRPS $\downarrow$ 0.391 $0.476$ 0.446 $0.462$ $0.471$ $0.474$ $0.466$ $0.655$ $0.713$ $0.710$ $0.666$ $1.000$
Networking MASE $\downarrow$ 0.635 $0.795$ $0.786$ $0.773$ $0.765$ $0.779$ 0.757 $0.871$ $0.856$ $0.933$ $1.035$ $1.000$
CRPS $\downarrow$ 0.400 $0.493$ 0.484 0.484 $0.493$ $0.506$ $0.489$ $0.725$ $0.721$ $0.739$ $0.734$ $1.000$
Security MASE $\downarrow$ 0.682 $0.741$ $0.739$ $0.736$ 0.717 $0.734$ $0.729$ $0.828$ $0.770$ $0.776$ $0.924$ $1.000$
CRPS $\downarrow$ 0.476 $0.505$ 0.504 0.504 $0.525$ $0.539$ $0.535$ $0.664$ $0.625$ $0.629$ $0.735$ $1.000$